OpenStudy (adamaero):

DE question: (2+m)^2 = 0

OpenStudy (adamaero):

So m_1 = -2 = m_2. But when I'm finding a differential equation to the auxillary equation, why is the "x" in there: y = c_1*e^(-2x) + c_2*x*e^(-2x) + ...

zepdrix (zepdrix):

Why is the general solution in the form of exponentials? Is that what you're asking?

OpenStudy (adamaero):

nope

zepdrix (zepdrix):

Oh oh the repeated solution! I see :)

zepdrix (zepdrix):

Hmm good question :3 I'm trying to remember why that is..

OpenStudy (adamaero):

I thought it was c_1*e^... (only)

OpenStudy (adamaero):

but the solution has a c_2*x*e^...

zepdrix (zepdrix):

http://tutorial.math.lamar.edu/Classes/DE/RepeatedRoots.aspx The top half of the page explains where that extra x is coming from with repeated roots. It's a little lengthy so I can't really comment on it. But the process makes sense.

OpenStudy (unklerhaukus):

\[(2+m)^2=0\\ m=-2,-2\\ y(x) = (A+Bx)e^{-2x}\]

OpenStudy (unklerhaukus):

repeated root ==> resonance